# PENEX: AdaBoost-Inspired Neural Network Regularization

### v2

@@ 6   Update sample weights:
Appendix C Practical Implementation of PENEX
C.1 Inference
Motivated by the Fisher consistency result (Prop. 2.1), we always rescale logits that are optimized on PENEX by $1+\alpha$. Specifically, to obtain a prediction $\hat{P}_{\scriptscriptstyle\textrm{PENEX}}$, we always use
@@ C.1 Inference
C.2 Default Value for Sensitivity $\alpha$
@@
If PENEX is used without tuning, we generally recommend 21 as the default choice. If, however, there is substantial label noise, we recommend a smaller value. If there is little label noise, $\alpha$ may be chosen larger, leading to more aggressive margin maximization.
C.3 Default Values for Alg. 1
@@
C.4 PyTorch-like Code
@@
Appendix D Experimental Details
D.1 Libraries
@@
D.2 Dataset & Model Licenses
@@ D.5 Metrics
D.6 Data Preprocessing
D.6.1 Computer Vision
@@
D.6.2 Fine-Tuning RoBERTa
@@
Table 2: Affects of the Sensitivity Parameter on Validation Curves. Smaller $\alpha$ typically lead to faster convergence, but worse generalization toward the end of the training duration (as can be seen more clearly for CIFAR-100). In addition, for very large $\alpha=3.2$, training curves tend to become less stable.

### v3

@@ 6   Update sample weights:
Appendix D Practical Implementation of PENEX
D.1 Inference
Motivated by Prop. 2.1, we rescale logits that are optimized on PENEX by $1+\alpha$. Specifically, to obtain a prediction $\hat{P}_{\scriptscriptstyle\textrm{PENEX}}$, we always use
@@ D.1 Inference
D.2 Default Value for Sensitivity $\alpha$
@@
If PENEX is used without tuning, we generally recommend 19 as the default choice. If, however, there is substantial label noise, we recommend a smaller value. If there is little label noise, $\alpha$ may be chosen larger, leading to more aggressive margin increase.
D.3 Default Values for Alg. 1
@@
D.4 PyTorch-like Code
@@
Appendix E Experimental Details
E.1 Libraries
@@
E.2 Dataset & Model Licenses
@@ E.5 Metrics
E.6 Data Preprocessing
E.6.1 Computer Vision
@@
E.6.2 Fine-Tuning RoBERTa
@@
Table 2: Effects of the Sensitivity Parameter on Validation Curves. Smaller $\alpha$ typically lead to faster convergence, but worse generalization toward the end of the training duration (as can be seen more clearly for CIFAR-100). In addition, for very large $\alpha=3.2$, training curves tend to become less stable.